OpenStudy (anonymous):

find the cos(3pi/2 -x) in terms of x

OpenStudy (anonymous):

-cosx

OpenStudy (anonymous):

how would i plug it into a calculator

OpenStudy (anonymous):

no bro

OpenStudy (anonymous):

no what

OpenStudy (raden):

it should be -sinx

OpenStudy (anonymous):

I DNT KNOW

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

my mistake cos(270- x)= -six

OpenStudy (anonymous):

whats six the number?

OpenStudy (raden):

sin(x), he means.. lol

OpenStudy (anonymous):

yeah bro

OpenStudy (anonymous):

my keyboard problem.

OpenStudy (anonymous):

im looking for x tho not sin is i the same thing

OpenStudy (anonymous):

is it the same thing

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how did u get 270

OpenStudy (anonymous):

180+90=270

OpenStudy (anonymous):

but theres a pi sogn in there like what did u do with that

OpenStudy (anonymous):

sign

OpenStudy (anonymous):

sin( 180+90)= -sin90= -1

OpenStudy (anonymous):

but what happened to the 3 and the pi

OpenStudy (anonymous):

sign depends on which quadrant angle belongs.

OpenStudy (anonymous):

ok so x is -1

OpenStudy (anonymous):

in first quadrant sin,cos,tan ,cot,cosec,sec all +ve,in second quadrant only sin,cosec +ve rest -ve,in 3rd quad only tan,cot +ve, rest -ve,in 4rd quadrant only cos,sec +ve,rest -ve.

OpenStudy (anonymous):

okay but is -1 in trems of x

OpenStudy (anonymous):

ok brother.r u clear.

OpenStudy (anonymous):

no but ill act like it

OpenStudy (anonymous):

pia in radian convert degree.